OpenStudy (anonymous):

how do you figure out in the 4th power?

OpenStudy (ghazi):

just multiply four times a^4= a*a*a*a

OpenStudy (anonymous):

so if its 3. i would do 3*3*3*3? so 3*3=9 and where do i go from there? do i multiply by 3 again or 9

OpenStudy (anonymous):

you multiply by 3 again 3*3*3*3 9*3*3 27*3 81

OpenStudy (anonymous):

thank you so much

hero (hero):

Yup, exponents is just shorthand for multiplication

hero (hero):

Just like multiplication is shorthand for addition.

OpenStudy (anonymous):

ok thank you. this is my first time doing exponents and class and thee are no examples or anything

hero (hero):

Shame on the teacher for not including examples. Lucky for you that you found OS

OpenStudy (phi):

you can do it a few ways, Here is another way 3*3*3*3 you can write as (3*3)*(3*3) or 9*9= 81

OpenStudy (anonymous):

how would i do 1.10 in the 5th power?

hero (hero):

(1.10)(1.10)(1.10)(1.10)(1.10) I think most of us would just use a calc, unless there's some shortcut available that I'm not aware of.

OpenStudy (phi):

if you have a scientific calculator, there is a button with ^ that you would use 1.1^5 you can also type 1.1^5= in the google search window